OpenStudy (anonymous):

Scoring a hole-in-one is the greatest shot a golfer can make. Once 8 professional golfers each made holes-in-one on the 6th hole at the same golf course at the same tournament. It has been found that the estimated probability of making a hole-in-one is 1/2153for male professionals. Suppose that a sample of 8 professional male golfers is randomly selected. What is the probability that at least one of these golfers makes a hole-in-one on the 15th hole at the same tournament?

OpenStudy (anonymous):

@sammixboo

sammixboo (sammixboo):

Ugh I hate probability give me a sec to try to figure this out.

sammixboo (sammixboo):

Sorry, can't help. I forgot how to do this

OpenStudy (anonymous):

It's fine thank you for trying !!

OpenStudy (kropot72):

Let the probability that none of the eight players makes a hole-in-one be P(0). Then the probability that at least one of these golfers makes a hole-in-one on the 15th hole at the same tournament is given by: 1.0000 - P(0) \[\large P(0)=(\frac{2152}{2153})^{8}\] Therefore \[\large P(none\ make\ a\ holeinone)=1.0000-(\frac{2152}{2153})^{8}=you\ can\ calculate\]